OpenStudy (anonymous):

evaluate ∫upper bound (3) lower bound (-2) (4t^3+6) dt

hartnn (hartnn):

you just need the x^n formula of integration \(\Large \int x^n dx = \dfrac{x^{n+1}}{n+1}+c\)

hartnn (hartnn):

for the first term 4 is constant for t^3, n = 3 for the 2nd term 6 is constant, t^0 so,n=0

OpenStudy (anonymous):

what i dont get is why the only possible answers are : -103 -95 95 103

OpenStudy (anonymous):

could you help me ??

hartnn (hartnn):

thats because after integration, we plug in the upper and lower limits... is this your first integration question ?? if not, have you tried to start to solve this one ?

OpenStudy (anonymous):

so wait, do you plug the numbers in ... (x^(3+1)/(3+1)) + 4 = x^4/4 + 4

hartnn (hartnn):

use 't' so for 4t^3 you get \(\Large 4\dfrac{t^4}{4}\) got this ? how about \(6=6t^0\) ?

OpenStudy (anonymous):

i dont get it .. how do you use t to get that equation?

hartnn (hartnn):

the variable in your question is 't' so, \(\Large \int t^n dt = \dfrac{t^{n+1}}{n+1}+c \\ \Large \int t^3 dt = \dfrac{t^{3+1}}{3+1}+c\) got it ?

OpenStudy (anonymous):

ooo yeah got it now !!

OpenStudy (anonymous):

so : \[6\frac{ t^0 }{ 1 }\] ?

hartnn (hartnn):

6t^1/1

hartnn (hartnn):

which is 6t

hartnn (hartnn):

\(\\ \Large \int 4t^3+6t^0 dt = 4\dfrac{t^{3+1}}{3+1}+6 \dfrac{t^1}{1}c = 4 \dfrac{t^4}{4}+6t+c\\ \Large = t^4+6t+c\)

hartnn (hartnn):

now substituting the limits (we don't consider the constant 'c') \(\Large [t^4+6t]^3_{-2} = ...\)

hartnn (hartnn):

first plug in the upper limit then lower limit and subtract \(\(\Large [t^4+6t]^3_{-2} = .[(3^4+6\times 3) - ((-2)^4+6\times (-2))] =...\)

hartnn (hartnn):

ignore that \( in the beginning

OpenStudy (anonymous):

95 !!!

OpenStudy (anonymous):

thank youuuuu !! i appreciate the help!!

hartnn (hartnn):

welcome ^_^